# Improving LLM First-Token Predictions in Multiple-Choice Question Answering via Prefilling Attack

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly evaluated on multiple-choice question answering (MCQA) tasks using First-Token Probability (FTP), which selects the answer option whose initial token has the highest likelihood. While efficient, FTP can be fragile: models may assign high probability to unrelated tokens (misalignment) or use a valid token merely as part of a generic preamble rather than as a clear answer choice (misinterpretation), undermining the reliability of symbolic evaluation. We propose a simple solution: the prefilling attack, a structured natural-language prefix (e.g., ‘The correct option is:’) prepended to the model output. Originally explored in AI safety, we repurpose prefilling to steer the model to respond with a clean, valid option, without modifying its parameters. Empirically, the FTP with prefilling strategy substantially improves accuracy, calibration, and output consistency across a broad set of LLMs and MCQA benchmarks. It outperforms standard FTP and often matches the performance of open-ended generation approaches that require full decoding and external classifiers, while being significantly more efficient. Our analysis suggests that prefilling is a simple, robust, and low-cost method to enhance the reliability of FTP-based evaluation in multiple-choice settings.
Improving LLM First-Token Predictions
in Multiple-Choice Question Answering via Prefilling Attack
Silvia Cappelletti1thanks: Equal contribution. Tobia Poppi1,211footnotemark: 1 Samuele Poppi1,2 Zheng-Xin Yong3 Diego Garcia-Olano4Marcella Cornia1Lorenzo Baraldi1Rita Cucchiara1 1University of Modena and Reggio Emilia  2University of Pisa 3 Brown University  4Meta {silvia.cappelletti, tobia.poppi, samuele.poppi, marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.it zheng_xin_yong@brown.edudiegoolano@meta.com
@@ 1 Introduction
Figure 1: We show that a simple prefilling attack template, which directs an LLM’s first generated token to a valid option for MCQA, substantially improves the standard first-token probability approach.
@@
Large Language Models (LLMs) are increasingly deployed as general-purpose reasoning systems (Plaat et al., 2024; Kojima et al., 2022; Huang and Chang, 2023): they are expected not only to generate fluent text, but also to make decisions (Liu et al., 2024; Lyu et al., 2025), answer questions (Liang et al., 2022; Kamalloo et al., 2023), and demonstrate understanding across a wide range of domains (Wei et al., 2022; Naveed et al., 2023). A common way to evaluate these capabilities is through question-answering tasks, where the model is required to process a question and return a correct and relevant answer (Liang et al., 2022; Kamalloo et al., 2023; Chang et al., 2024; Khashabi et al., 2020). Among the many QA formats, one of the most widely adopted is multiple-choice question answering (MCQA), in which the model selects from a fixed set of answer candidates, typically labeled A through D (Hendrycks et al., 2021; Lai et al., 2017; Clark et al., 2018; Richardson et al., 2013; Li et al., 2024).
MCQA benchmarks such as MMLU (Hendrycks et al., 2021) are widely used to evaluate the general knowledge, reasoning ability, and decision-making skills of a model across diverse subjects. Models are typically prompted with a question and a set of answer options, and evaluation is done by checking the model choice across the given, valid options.
@@
However, there are various ways with differing tradeoffs to extract the model choice. Some approaches decode free-form generations into valid options either by using a separate language model (e.g., GPT-3.5-Turbo Ouyang et al. (2022)) or by training a dedicated classifier (Wang et al., 2024a). Another approach that does not require an additional model for extraction is to let the model generate one single token, and compare its probability across the set of valid option tokens (e.g., ‘A’, ‘B’, ‘C’, ‘D’). This method, known as First-Token Probability (FTP) Hendrycks et al. (2021); Santurkar et al. (2023), scores each candidate answer by computing the likelihood of it being generated as the first token and selecting the most probable one, without requiring full autoregressive decoding. It is efficient and model-agnostic, but relies on the assumption that the model is ready to commit to an answer immediately – an assumption that might break down in practice.
@@
In particular, models frequently diverge from the expected output structure, producing full sentences (e.g., ‘I believe the answer is A’), or ambiguous strings that are hard to interpret (i.e., first-token misalignment). Worse still, we find that even when the first token is a valid label, it may serve a purely grammatical role that does not reflect the model intended answer (i.e., first-token misinterpretation). For instance, ‘A’ could begin ‘A possible answer could be C’. These ambiguities make first-token evaluation noisy and potentially misleading (Figure 1), eventually distorting accuracy metrics.
@@
In this work, we introduce a simple, yet effective, solution to these two problems by using the prefilling attack Tang (2024); Andriushchenko et al. (2025) to guide the LLM towards generating a valid option as the first token. Specifically, we draw on the underlying principle of biasing model behavior through structured prefilling and show that by prepending a short and benign phrase (such as ‘The correct option is: ’) to the model generation, it is possible to effectively steer the model toward generating a valid multiple-choice response. We show that our approach leads to consistent improvements across different LLMs and benchmarks (e.g., up to +$40$% on the Gemma-2-9B model (Team et al., 2024), when tested on the MMLU benchmark).
@@
Despite its simplicity, this prefilling strategy substantially improves the reliability of FTP-based evaluation by reducing both the first-token misalignment and first-token misinterpretation issues. To further validate this, we conduct controlled experiments where the model selects its first token from the full vocabulary, rather than a filtered set as in standard FTP, allowing us to observe whether the top-ranked token aligns with the intended answer. Prefilling consistently steers the model toward ranking the correct option highest, substantially reducing first-token misalignment – for instance, boosting Llama-3.1-8B’s accuracy on SciQ (Welbl et al., 2017) from 2.2% to 96.9%. Compared to stronger baselines (e.g., open-ended generation with GPT-based answer extraction), prefilling yields similarly aligned answers while being more efficient, as it avoids relying on an additional model for evaluation. For example, with Phi-4-14B, the alignment gap with this baseline drops from 7.6% to 1.6%, on the OpenBookQA dataset (Mihaylov et al., 2018).
To better understand the broader impact of prefilling beyond accuracy, we also examine its effect on model calibration (i.e., the ability to reflect uncertainty in predictions). For this purpose, we also evaluate whether prefilling improves the model confidence estimates. We find that prefilling consistently improves output calibration, lowering the expected calibration error across all tested models.
Overall, our analysis demonstrates the effectiveness of the prefilling attack as a low-cost and easily applicable intervention for improving the reliability of LLM outputs in multiple-choice settings.
@@ 2 Related Work
LLMs as Classifiers. Large Language Models (LLMs) have been increasingly used as zero-shot and few-shot classifiers by leveraging their in-context learning abilities Brown et al. (2020); Chowdhery et al. (2023); Achiam et al. (2023). In multiple-choice question answering (MCQA), LLMs are typically prompted with a question followed by a list of labeled answer candidates. The model is then expected to produce or select the correct option, either via direct generation or scoring. Recent benchmarks such as MMLU Hendrycks et al. (2021), RACE Lai et al. (2017), and ARC Clark et al. (2018) have become standard tools for assessing LLM performance across a range of domains.
@@
To operationalize classification within a generative framework, these models are typically evaluated by measuring how likely they are to produce the correct label at the beginning of their output (Holtzman et al., 2021; Zhang et al., 2022; Min et al., 2022). This requires mapping free-form text generation into a constrained prediction task – often by prompting the model to respond with a symbolic label such as ‘A’ or ‘B’, corresponding to multiple-choice answers. Such symbolic setups allow for structured evaluation without retraining, but also introduce new sources of fragility tied to the surface-level generation behavior of the model (Wang et al., 2024a, b).
@@
First-Token Probability and Variants. A common method for MCQA with LLMs is to decode the model output and evaluate whether the first generated token corresponds to the correct answer label Zhang et al. (2022). This first-token prediction approach is appealing due to its simplicity and compatibility with generative models. However, it is also sensitive to decoding randomness, tokenization artifacts, and inconsistent formatting, often leading to ambiguous or incorrect outputs Holtzman et al. (2021). Variants of this approach include ranking candidate completions by log-likelihood Brown et al. (2020); Zhang et al. (2022), constraining decoding to answer tokens only Zhao et al. (2021); Holtzman et al. (2021), or scoring options directly with logprob-based classification Min et al. (2022); Yao et al. (2023).
@@
Prefilling Attack and Prompt Injection. In the domain of AI safety, the prefilling attack Tang (2024); Andriushchenko et al. (2025) has emerged as a simple yet powerful prompt injection technique. It involves inserting innocuous-looking natural-language phrases into the model’s output prompt to subvert safety filters or steer the model toward undesired behavior Zou et al. (2023); Wei et al. (2023). For instance, adding a phrase such as ‘Sure! The answer is:’ can cause the model to comply with otherwise restricted user queries. These techniques expose vulnerabilities in even strongly aligned models, revealing the brittleness of instruction-following and safety mechanisms. While our work is not focused on bypassing safety constraints, we adopt the prefilling mechanism as a means of controlled behavioral biasing in MCQA, demonstrating its unintended but beneficial effects on model performance.
3 First-Token Probability and Prefilling Attack
@@ 3.1 First-Token Probability
MCQA and First-Token Probability. In MCQA, a model receives as input a question $q$ and a fixed set of symbolic answer options $\mathcal{A}^{q}=\{\text{`{A}'},\text{`{B}'},\text{`{C}'},\text{`{D}'}\}$ ’ }, and must select the most appropriate answer among them. To simplify evaluation, a common strategy is to restrict the probability distribution of the model to this small set of symbolic tokens, allowing for discrete scoring methods. FTP evaluates the model based solely on the probability it assigns to each answer token as the first word in its generated response. Given a formatted prompt (e.g., ‘Respond with the correct option: What is the capital of Italy? Answer:’), the model prediction is computed as
$\arg\max_{i}P(a^{q}_{i}),\leavevmode\nobreak\ \leavevmode\nobreak\ a^{q}_{i}\in\mathcal{A}^{q}$ (2)
@@
First-Token Misalignment and Misinterpretation. While the first-token probability approach is efficient and simple to implement, it suffers from two major failure modes. First, in first-token misalignment, the model may diverge from the expected symbolic output format, instead generating full sentences (e.g., ‘I believe the answer is A’) or ambiguous strings that begin with irrelevant tokens. In such cases, the correct answer may be present later in the sequence, but the real, predicted token lies outside the valid answer set, making evaluation misleading. Second, even when the model first token is a valid label (e.g., ‘A’), it may serve a purely grammatical or stylistic function rather than indicating the intended answer—we call this first-token misinterpretation. For example, a response like ‘A possible answer could be C’ begins with a token in the answer set (‘A’) but ultimately points to a different choice (‘C’), making the intended meaning unclear. These two phenomena introduce noise in model predictions, distort accuracy metrics, and hinder fair comparisons across prompting strategies and model architectures.
@@ 3.2 Prefilling Strategy
From AI Safety to MCQA. To address the limitations of standard FTP evaluation, we propose a constructive adaptation of the prefilling attack, a technique originally introduced in the AI safety community Tang (2024); Andriushchenko et al. (2025). In its adversarial form, the prefilling attack involves injecting a seemingly innocuous natural-language prefix (e.g., ‘Sure! Here’s the answer:’) to manipulate the model into generating otherwise restricted or unsafe content, effectively bypassing content filters. We repurpose this idea by prepending a short, benign prefix (e.g., ‘The correct option is:’) to the model’s response prompt, to steer the model toward producing a valid symbolic answer as its very first token. This intervention reduces ambiguity and misalignment by reinforcing the expected output format and encouraging direct, structured responses in MCQA tasks.
Combining FTP and Prefilling. Under our approach, the FTP method remains unchanged conceptually but is now applied to the modified prompt, which includes the prefilling prefix. This combined FTP-prefilling method preserves FTP efficiency and simplicity, while substantially mitigating issues of misalignment and misinterpretation, resulting in improved reliability, accuracy, and calibration across MCQA benchmarks.
In practice, to effectively inject these fixed tokens into the model output, we leverage the structured prompting conventions utilized by instruction-tuned LLMs. These models typically adhere to a specific dialogue format delineated by special control tokens (e.g., ‘<|user|>’ and ‘<|assistant|>’), defining the response structure clearly. Although the exact formatting varies between models–for example, ChatML used in OpenAI models, Alpaca-style formatting for Llama derivatives, or ChatGLM native dialog structure–the prefilling prefix can always be placed immediately following the assistant starting token, effectively conditioning the model’s subsequent generation on the injected tokens.
Figure 2: Visual examples of our prefilling strategy.
Our key insight is that by inserting the prefilling prefix directly into the assistant output turn (Figure 2), the model treats this injected text as previously generated content, naturally continuing its response from the provided context and thus aligning closely with the expected MCQA format.

### v2

@@
11institutetext: University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.it
22institutetext: University of Pisa, Italy33institutetext: Mohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.ae
44institutetext: Brown University, United States
44email: zheng_xin_yong@brown.edu
55institutetext: Meta, United States
55email: diegoolano@meta.com
Improving LLM First-Token Predictions in Multiple-Choice Question Answering via
Output Prefilling
Silvia Cappelletti1 The first two authors equally contributed to this research.University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
[…]
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
Samuele Poppi3
This work was carried out primarily while the author was affiliated with the University of Modena and Reggio Emilia, Italy.University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
   Zheng Xin Yong4 University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
   Diego Garcia-Olano5 University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
Marcella Cornia1
University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
   Lorenzo Baraldi1 University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
   Rita Cucchiara1 University of Modena and Reggio Emilia, Italy
11email: {silvia.cappelletti, tobia.poppi,
11email: marcella.cornia, lorenzo.baraldi, rita.cucchiara}@unimore.itUniversity of Pisa, ItalyMohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
33email: samuele.poppi@mbzuai.ac.aeBrown University, United States
44email: zheng_xin_yong@brown.eduMeta, United States
55email: diegoolano@meta.com
@@ Abstract
Large Language Models (LLMs) are traditionally evaluated on multiple-choice question answering (MCQA) tasks using First-Token Probability (FTP), which selects the answer option whose initial token has the highest likelihood. While efficient, FTP can be fragile: models may assign high probability to unrelated tokens (misalignment) or use a valid token merely as part of a generic preamble rather than as a clear answer choice (misinterpretation), undermining the reliability of symbolic evaluation. We propose a simple solution: output prefilling, a structured natural-language prefix (e.g., ‘The correct option is:’) prepended to the model output. Originally explored in AI safety as an attack strategy, we repurpose prefilling to steer the model to respond with a clean, valid option, without modifying its parameters. Through extensive evaluation, we find that the FTP with prefilling strategy substantially improves accuracy, calibration, and output consistency across a broad set of LLMs and MCQA benchmarks. It outperforms standard FTP and often matches the performance of open-ended generation approaches that require full decoding and external classifiers, while being significantly more efficient. Our analysis suggests that prefilling is a simple, robust, and zero-cost method to enhance the reliability of FTP-based evaluation in multiple-choice settings.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly deployed as general-purpose reasoning systems [15, 17, 33], where they are expected not only to generate fluent text, but also to make decisions [23, 24], answer questions [21, 16], and demonstrate understanding across a wide range of domains [42, 30]. A common way to evaluate these capabilities is through question-answering tasks, where the model is required to process a question and return a correct and relevant answer [21, 16, 6]. Among the many QA formats, one of the most widely adopted is multiple-choice question answering (MCQA), in which the model selects from a fixed set of answer candidates, typically labeled A through D [13, 18, 8, 20]. MCQA benchmarks such as MMLU [13] are widely used to evaluate the general knowledge, reasoning ability, and decision-making skills of a model across a broad range of subjects. In practice, models are typically prompted with a question and a set of answer options, and evaluation is performed by checking the model choice across the given, valid options.
@@
There are various ways with differing tradeoffs to extract the model choice. Some approaches decode free-form generations into valid options either by using an auxiliary language model (e.g., GPT-3.5-Turbo [32]) or a dedicated classifier [46, 39]. While constraining outputs to a fixed format via an external model leads to more reliable evaluation, free-form generation alone may reduce alignment with human judgment [27]. Another approach, which does not require an external model, is to let the model generate one single token, and compare its probability across the set of valid option tokens only (e.g., ‘A’, ‘B’, ‘C’, ‘D’). This method, known as First-Token Probability (FTP) [13, 35], scores each candidate answer by computing the likelihood of it being generated as the first token and selecting the most probable one, without requiring full autoregressive decoding.
@@
While FTP is efficient and model-agnostic, it relies on the assumption that the model is ready to commit to an answer immediately – an assumption that might break down in practice. In particular, models frequently diverge from the expected output structure, producing full sentences (e.g., ‘I believe the answer is A’), or ambiguous strings that are hard to interpret (i.e., first-token misalignment). Worse still, we find that even when the first token is a valid label, it may serve a purely grammatical role that does not reflect the model intended answer (we call this first-token misinterpretation). For instance, ‘A’ could begin ‘A possible answer could be C’. These ambiguities make first-token evaluation noisy and potentially misleading (Fig. 1), eventually distorting accuracy metrics.
Figure 1: We show that a simple output prefilling template, which directs an LLM’s first generated token to a valid option for MCQA, substantially improves the standard first-token probability approach.
To make things worse, prior work has shown that MCQA model performance is highly sensitive to prompt design. Small variations in choice symbols, choice ordering, or phrasing can substantially affect model outputs, and thus their evaluation scores [27, 4]. This motivates why holding inputs fixed is needed for established benchmarks to ensure fair comparisons, and why relying on prompt design alone can not fully guarantee correct first-token outputs in zero-shot FTP evaluation, leaving a persistent risk of malformed or misaligned responses.
@@
To address these limitations, in this work we introduce a simple yet effective solution by repurposing the prefilling attack, originally introduced in AI safety [38, 3], to guide LLMs towards generating a valid option as the first token. Specifically, we build on the principle of biasing model behavior through structured prefilling and show that by prepending a short and benign phrase (such as ‘The correct option is: ’) effectively steers the model toward generating valid multiple-choice responses. Our experiments demonstrate that, despite its simplicity, this output prefilling approach substantially improves the reliability of FTP-based evaluation by reducing both the first-token misalignment and first-token misinterpretation issues, thus leading to consistent accuracy gains across different LLMs and benchmarks.
@@
To further validate this, we conduct controlled experiments in which the model selects its first token from the full vocabulary, rather than from the filtered valid options used in standard FTP, enabling a direct test of whether the top-ranked token matches the intended answer. Across models, prefilling consistently steers predictions toward the correct option, substantially reducing first-token misalignment. When compared against stronger baselines (e.g., open-ended generation with GPT-based answer extraction), prefilling yields similarly aligned answers while being more efficient, as it removes the need for an auxiliary evaluation model. Beyond accuracy, we also assess its impact on model calibration (i.e., the ability to represent predictive uncertainty) and observe consistent, substantial improvements across all models and settings, indicating that prefilling enhances both performance and robustness.
Overall, our analysis shows output prefilling to be a simple zero-cost intervention that substantially improves the reliability of LLM outputs across diverse MCQA benchmarks. To our knowledge, this is the first work to rigorously quantify its effectiveness with modern general-purpose LLMs.
@@ 2 Related Work
LLMs as Classifiers. LLMs have been increasingly used as zero-shot and few-shot classifiers by leveraging their in-context learning abilities [5, 7, 1]. In MCQA, LLMs are typically prompted with a question followed by a list of labeled answer candidates. The model is then expected to produce or select the correct option, either via direct generation or scoring. Recent benchmarks such as MMLU [13], RACE [18], and ARC [8] have become standard tools for assessing LLM performance across a range of domains. To operationalize classification within a generative framework, models are typically evaluated by measuring how likely they are to produce the correct label at the beginning of their output [14, 48, 26]. This requires mapping free-form text generation into a constrained prediction task, typically by prompting the model to output a symbolic label (e.g., ‘A’ or ‘B’) corresponding to multiple-choice answers. Such symbolic setups allow for structured evaluation without retraining, but also introduce fragility tied to the surface-level generation behavior of the model [39, 40].
@@
First-Token Probability and Variants. A common method for MCQA with LLMs is to decode the model output and evaluate whether the first generated token corresponds to the correct answer label [48]. This first-token prediction approach is appealing due to its simplicity and compatibility with generative models. However, it is also sensitive to decoding randomness, tokenization artifacts, and inconsistent formatting, often leading to ambiguous or incorrect outputs [14]. Variants of this approach include ranking candidate completions by log-likelihood [5, 48], constraining decoding to answer tokens only [49, 14], or scoring options directly with logprob-based classification [26, 45].
@@
Prefilling Attack and Prompt Injection. In AI safety, the prefilling attack [38, 3] has emerged as a simple yet powerful prompt injection technique. It involves inserting innocuous-looking natural-language phrases into the model’s output prompt to subvert safety filters or steer the model toward undesired behavior [51, 41]. For instance, adding a phrase such as ‘Sure! The answer is:’ can cause the model to comply with otherwise restricted user queries. These techniques expose vulnerabilities in even strongly aligned models, revealing the brittleness of instruction-following and safety mechanisms. While our work is not focused on bypassing safety constraints, we adopt prefilling as a means of controlled behavioral biasing in MCQA, demonstrating its unintended but beneficial effects on model performance. While techniques like chain-of-thought [43], self-ask [34], and tree-of-thoughts [45] involve guiding the model reasoning process through structured prompts, they do not guarantee strict adherence to a specific output format. In contrast, output-side prefilling directly inserts a fixed prefix into the model output, exploiting the normal cognitive biases of the model and thus ensuring that the first generated token aligns with the desired answer. This method enforces format consistency mechanically, which is particularly beneficial for tasks like FTP evaluation, where precise output formatting is crucial.
3 First-Token Probability and Output Prefilling
@@ 3.1 First-Token Probability
MCQA and First-Token Probability. In MCQA, a model receives as input a question $q$ and a fixed set of symbolic answer options $\mathcal{A}^{q}=\{\text{`{\small A}'},\text{`{\small B}'},\text{`{\small C}'},\text{`{\small D}'}\}$, and must select the most appropriate answer among them. To simplify evaluation, a common strategy is to restrict the probability distribution of the model to this small set of symbolic tokens, allowing for discrete scoring methods. FTP evaluates the model based solely on the probability it assigns to each answer token as the first word in its generated response. Given a formatted prompt (e.g., ‘Respond with the correct option: What is the capital of Italy? Answer:’), the model prediction is computed as
$\arg\max_{i}P(a^{q}_{i}),~~a^{q}_{i}\in\mathcal{A}^{q}$ (2)
@@
First-Token Misalignment and Misinterpretation. While the first-token probability approach is efficient and simple to implement, it suffers from two major failure modes. First, in first-token misalignment, the model may diverge from the expected symbolic output format, instead generating full sentences (e.g., ‘I believe the answer is A’) or ambiguous strings that begin with irrelevant tokens. In such cases, the correct answer may be present later in the sequence, but the real, predicted token lies outside the valid answer set, making evaluation misleading. Second, even when the model first token is a valid label (e.g., ‘A’), it may serve a purely grammatical or stylistic function rather than indicating the intended answer – we call this first-token misinterpretation. For example, a response like ‘A possible answer could be C’ begins with a token in the answer set (‘A’) but ultimately points to a different choice (‘C’), making the intended meaning unclear. These two phenomena introduce noise in model predictions, distort accuracy metrics, and hinder fair comparisons across prompting strategies and model architectures.
@@ 3.2 Prefilling Strategy
From AI Safety to MCQA. To address the limitations of standard FTP evaluation, we propose an adaptation of the prefilling attack, a technique originally introduced in the AI safety community [38, 3]. In its adversarial form, the prefilling attack involves injecting a seemingly innocuous natural-language prefix (e.g., ‘Sure! Here’s the answer:’) to manipulate the model into generating otherwise restricted or unsafe content, effectively bypassing content filters. We repurpose this idea by prepending a short, benign prefix (e.g., ‘The correct option is:’) to the model response prompt, to steer the model toward producing a valid symbolic answer as its very first token. This output prefilling intervention reduces ambiguity and misalignment by reinforcing the expected output format and encouraging direct, structured responses in MCQA tasks.
Figure 2: Visual examples of our prefilling strategy. The prefilling template is added after the assistant’s response start to mechanically condition the output format.
Combining FTP and Prefilling. Under our approach, the FTP method remains unchanged conceptually but is now applied after the assistant control token together with our prefilling prefix. This combined FTP–prefilling method preserves the efficiency and simplicity of FTP, while substantially mitigating issues of misalignment and misinterpretation, resulting in improved reliability, accuracy, and calibration across MCQA benchmarks.
In practice, to effectively inject these fixed tokens into the model output, we leverage the structured prompting conventions utilized by instruction-tuned LLMs. These models typically adhere to a specific dialogue format delineated by special control tokens (e.g., ‘<|user|>’ and ‘<|assistant|>’), defining the response structure clearly. Although the exact formatting varies between models (for example, ChatML used in OpenAI models, Alpaca-style formatting for Llama derivatives, or ChatGLM native dialog structure), the prefilling prefix can always be placed immediately following the assistant starting token, effectively conditioning the subsequent generation of the model on the injected tokens.
Our key insight is that by inserting the prefilling prefix directly into the assistant output turn (Fig. 2), the model treats this injected text as previously generated content, naturally continuing its response from the provided context and thus aligning closely with the expected MCQA format.